Prove that the ratio of magnetic energy density and electric energy density is unity.
1
Expert's answer
2020-04-23T12:48:33-0400

In an electromagnetic wave

"E=\\sqrt{\\dfrac{\u200b\u03bc_o}{\\epsilon_o}}\u200b\u200b\n\u200bH" ..................(1)

Magnetic Field energy density :

"η_B=\dfrac{H^2}{2μ_o}"

Electric Field energy density :

"η_E=\dfrac{E^2ϵ_o}{2}"

Substituting the value of E in the in the ηE we get "\u03b7_E=\u03b7_B"